# Continuum limit of the mobility edge and taste-degeneracy effects in high-temperature lattice QCD with staggered quarks

### 05 Dec 2023 v1

@@ 3.1 Lattice setup
In this work we considered 5 ensembles of gauge configurations of $N_{f}=2+1$ QCD at the physical point originally generated for the investigation of the QCD topological susceptibility at finite temperature reported in Ref. [66], obtained on $N_{s}^{3}\times N_{t}$ lattices with lattice spacings $a$ ranging from $\sim 0.107\,\mathrm{fm}$ down to $\sim 0.054\,\mathrm{fm}$. In all cases, the temperature was fixed to $T=1/(aN_{t})\simeq 230\,\mathrm{MeV}$, and the lattice size to $L=aN_{s}\simeq 3.4\,\mathrm{fm}$, corresponding to an aspect ratio $N_{s}/N_{t}=LT=4$. For the next-to-finest lattice spacing, corresponding to $N_{t}=14$, we also considered 2 additional ensembles with different lattice volumes, with aspect ratios $LT\simeq 3.4$ and $\simeq 4.6$, corresponding respectively to $L\simeq 2.9\,\mathrm{fm}$ and $\simeq 3.9\,\mathrm{fm}$. In order to keep our paper self-contained, we briefly summarize below the lattice setup employed in Ref. [66] to generate configurations, and we refer the reader to the original reference for further technical details.
@@
The Monte Carlo simulations of Ref. [66] employed Rational Hybrid Monte Carlo (RHMC) updating algorithm [75, 76, 77], used, for some simulation points, in combination with the multicanonical algorithm [78, 79, 80, 81, 66, 82]. As a matter of fact, it is well known that, above the QCD chiral crossover $T_{c}\simeq 155\,\mathrm{MeV}$, the topological susceptibility $\chi=\langle Q^{2}\rangle/V_{4}$, where $V_{4}=L^{3}/T$, is rapidly suppressed as a function of the temperature [83, 84, 85, 66], leading to $\langle Q^{2}\rangle=V_{4}\chi\ll 1$. Multicanonical simulations are employed to enhance the probability of visiting nontrivial topological sectors when this probability is suppressed, without spoiling importance sampling. For what concerns the present investigation, the multicanonical algorithm allows one to avoid possible systematic effects in the low-lying staggered spectrum, which is connected to topological excitations. Concerning the technical details of multicanonical runs, we refer the reader to the original work [66].
@@
Table 1: Simulation parameters used in [66] to generate the physical-point gauge ensembles considered in this work. The bare parameters $\beta$, $am_{s}$, $am_{l}=am_{s}/28.15$ and the lattice spacing $a$ have been fixed according to the results of Refs. [72, 73, 74], and correspond to physical pion mass and physical strange-to-light quark mass ratio. Simulations marked with * have been performed without the multicanonical algorithm.
@@
For each gauge configuration we then analyzed the lowest 150 positive eigenvalues of $D^{\mathrm{stag}}[U^{(2)}]$ [i.e., the same operator appearing in the fermionic determinant, Eq. (17)], obtained in Ref. [66] using the PARPACK library [86], from which we computed the corresponding unfolded eigenvalues. We complemented these data by performing a new calculation to obtain the IPRs of the Dirac eigenvectors for the gauge ensembles with $N_{t}=14$. An example of a staggered Dirac spectrum on a single, typical configuration is shown in Fig. 1. We performed unfolding by ordering all the available eigenvalues $\lambda_{n}$ of all the configurations of a given lattice ensemble, and replacing them with their rank divided by the number of configurations [17]. This automatically yields unit spectral density, while the deviation of $\langle s\rangle$ from 1 can be used to test the accuracy of the procedure.
@@
Figure 2: Top panel: comparison of $\langle s\rangle$ as a function of $\lambda/m_{s}$ for all the explored lattice spacings at fixed value of the aspect ratio $LT=4$. Bottom panel: comparison of $\langle s\rangle$ as a function of $\lambda/m_{s}$ for all the explored aspect ratios, corresponding to different values of $N_{s}$, at fixed value of the lattice spacing, corresponding to a temporal extent $N_{t}=14$. Our continuum estimate for the mobility edge (see below Sec. 3.5) is marked in both panels by a red circle on the horizontal axis.
@@ 3.3 Determination of the mobility edge from spectral statistics
Figure 4: Comparison of $I_{s_{0}}$ (top panel) and $\mathrm{Var}(s)$ (bottom panel), computed locally in the spectrum on lattices at fixed temperature $T=230\,\mathrm{MeV}$, for different values of the lattice spacing at fixed spatial lattice size $L=3.4\,\mathrm{fm}$ in physical units (i.e., fixed $LT=4$). Our continuum estimate for the mobility edge (see below Sec. 3.5) is marked by a red circle on the horizontal axis.
Figure 5: Comparison of $I_{s_{0}}$ (top panel) and $\mathrm{Var}(s)$ (bottom panel), computed locally in the spectrum on lattices at fixed temperature $T=230\,\mathrm{MeV}$, for different spatial volumes at fixed lattice spacing, $a=0.0615\,\mathrm{fm}$ (i.e., $N_{t}=14$). Our continuum estimate for the mobility edge (see below Sec. 3.5) is marked by a red circle on the horizontal axis.
@@
The data show that the two coarsest lattices ($N_{t}=8,10$) give compatible results for the spectral statistics in all spectral regions in the bulk of the spectrum, and down to the first bin below the mobility edge. Barring a conspiracy, this indicates two things: that the formation of taste multiplets does not have significant effects in that spectral region yet; and that, when this is the case, further lattice artifacts, on top of those introduced by taste-degenerate multiplets, are small when considering localization properties at fixed physical volume. This shows that the determination of the mobility edge from the critical spectral statistics is reliable on these lattices. Clearly, it may still be affected by finite-volume systematics, but these would involve only effects unrelated to taste multiplets (such as the localization length in certain spectral regions near the mobility edge being too large compared with the spatial size of the system), which are expected not to affect much the determination of the mobility edge via the matching with the critical statistics (see above in Sec. 2.3).
The dependence of $\lambda_{c}^{\mathrm{stat}}/m_{s}$ on the aspect ratio at fixed lattice spacing is visible instead in Fig. 5, where $N_{t}=14$ and the spatial size is varied. Here a larger aspect ratio drives $I_{s_{0}}$ and $\mathrm{Var}(s)$ down, as expected, so leading to an estimate for the mobility edge that decreases with the spatial volume. Notice that the scale-invariant nature of the mobility edge is still masked here by the distortions of the unfolded level spacing distribution, and no volume-independent crossing point of the various curves is present. This shows that the effects of taste degeneracy are still strong even on our largest lattice at $N_{t}=14$. This is in agreement with the argument discussed in Sec. 2.4, as, according to Eq. (15), we would need an aspect ratio of $LT\sim 9$ at this lattice spacing in order to have the same $\delta/\Delta$ that we have for $N_{t}=8$ with $LT=4$.
@@ 3.4 Taste-degeneracy effects on the mobility edge in the thermodynamic limit
In order to show that the effects of taste degeneracy become irrelevant for the determination of the mobility edge with our method when taking the thermodynamic limit at fixed lattice spacing, we have compared the $N_{s}\to\infty$ extrapolation of our $N_{t}=14$ data with an independent determination of $\lambda_{c}$ based directly on the properties of the eigenvectors. The infinite-volume extrapolation is shown in Fig. 7; the corresponding results for the mobility edge are:
@@
This is in good agreement with the thermodynamic extrapolation of our estimates from spectral statistics, Eq. (20), see Fig. 7. This shows that these are indeed converging to the actual position of the mobility edge. Moreover, this further confirms that finite-size effects on the estimate based on the critical fractal dimension are reasonably small.
@@
3.5 Dependence of the mobility edge on the lattice spacing
@@
We finally summarize the results of the previous two subsections in Fig. 10, showing together all our determinations of $\lambda_{c}/m_{s}$. It is clear that our two coarsest lattice give compatible estimates via $\lambda_{c}^{\mathrm{stat}}/m_{s}$, while on finer lattices with the same aspect ratio these estimates rapidly deviate. On the other hand, the infinite-volume extrapolation at $N_{t}=14$ gives again results compatible with our coarsest lattices, as well as with the determination from the local fractal dimension.
@@
Before attempting an extrapolation of $\lambda_{c}/m_{s}$ to the continuum, it is important to discuss whether the order in which one takes the thermodynamic and continuum limits matters. In this context, our estimate based on $\alpha(\lambda)$, being unaffected by taste degeneracy, allows us to define a valid estimate of the position of the mobility edge at both finite aspect ratio and finite spacing. Our numerical results show that this quantity does not depend strongly on the aspect ratios employed. Moreover, our results obtained at $N_{t}=14$ are not very different from the estimates obtained using spectral statistics on our coarsest lattices $N_{t}=8,10$, where taste-degeneracy effects are negligible and estimates based on critical statistics should accurately reflect the position of the mobility edge in the thermodynamic limit. These observations suggest that the estimate of the mobility edge based on the fractal dimension does not depend strongly neither on the aspect ratios nor on the lattice spacing. There is therefore no reason to expect the continuum and thermodynamic limit not to commute, when measuring the position of the mobility edge directly from the eigenvectors. Thus, since the thermodynamic limit of this estimate at fixed spacing matches that of the estimate from spectral statistics, we expect that taking the thermodynamic limit followed by the continuum limit of the latter provides the correct value of the mobility edge in the continuum theory.
@@
In the light of this discussion, we have performed an extrapolation to the continuum limit using our estimates of $\lambda_{c}/m_{s}$ from spectral statistics for $N_{t}=8,10$ and $N_{t}=14$, where for $N_{t}=14$ we have used the results extrapolated first towards the thermodynamic limit. The results obtained from the two observables $I_{s_{0}}$ and $\mathrm{Var}(s)$ are:
$\frac{\lambda_{c}}{m_{s}}$ $=1.748(75),\qquad$ (from $(\text{from }I_{s_{0}}),$ (24)
@@
$\frac{\lambda_{c}}{m_{s}}$ $=1.816(43),\qquad$ (from $(\text{from }\mathrm{Var}(s)),$ (25)
@@
perfectly compatible within errors. To our knowledge, this result is the first fully controlled extrapolation of $\lambda_{c}/m_{s}$ to the continuum.

### 12 Mar 2024 v2

@@ 3.1 Lattice setup
In this work we considered 5 ensembles of gauge configurations of $N_{f}=2+1$ QCD at the physical point, originally generated for the investigation of the QCD topological susceptibility at finite temperature reported in Ref. [75], obtained on $N_{s}^{3}\times N_{t}$ lattices with lattice spacings $a$ ranging from $\sim 0.107\,\mathrm{fm}$ down to $\sim 0.054\,\mathrm{fm}$. In all cases, the temperature was fixed to $T=1/(aN_{t})\simeq 230\,\mathrm{MeV}$, and the lattice size to $L=aN_{s}\simeq 3.4\,\mathrm{fm}$, corresponding to an aspect ratio $N_{s}/N_{t}=LT=4$. For the next-to-finest lattice spacing, corresponding to $N_{t}=14$, we also considered 2 additional ensembles with different lattice volumes, with aspect ratios $LT\simeq 3.4$ and $\simeq 4.6$, corresponding respectively to $L\simeq 2.9\,\mathrm{fm}$ and $\simeq 3.9\,\mathrm{fm}$. In order to keep our paper self-contained, we briefly summarize below the lattice setup employed in Ref. [75] to generate configurations, and we refer the reader to the original reference for further technical details.
@@
The Monte Carlo simulations of Ref. [75] employed the Rational Hybrid Monte Carlo (RHMC) updating algorithm [84, 85, 86], used, for some simulation points, in combination with the multicanonical algorithm [87, 88, 89, 90, 75, 91]. As a matter of fact, it is well known that, above the QCD chiral crossover $T_{c}\simeq 155\,\mathrm{MeV}$, the topological susceptibility $\chi=\langle Q^{2}\rangle/V_{4}$, where $V_{4}=L^{3}/T$, is rapidly suppressed as a function of the temperature [92, 93, 94, 95, 96, 97, 75], leading to $\langle Q^{2}\rangle=V_{4}\chi\ll 1$ for the lattices employed in typical simulations. Multicanonical simulations are employed to enhance the probability of visiting nontrivial topological sectors when this probability is suppressed, without spoiling importance sampling. For what concerns the present investigation, the multicanonical algorithm allows one to avoid possible systematic effects in the low-lying staggered spectrum, which is connected to topological excitations. Concerning the technical details of multicanonical runs, we refer the reader to the original work [75].
@@
Table 1: Simulation parameters used in Ref. [75] to generate the physical-point gauge ensembles considered in this work. The bare parameters $\beta$, $am_{s}$, $am_{l}=am_{s}/28.15$ and the lattice spacing $a$ have been fixed according to the results of Refs. [81, 82, 83], and correspond to physical pion mass and physical strange-to-light quark mass ratio. Simulations marked with * have been performed without the multicanonical algorithm.
@@
For each gauge configuration we then analyzed the lowest 150 positive eigenvalues of $D^{\mathrm{stag}}[U^{(2)}]$ [i.e., the same operator appearing in the fermionic determinant, Eq. (34)], obtained in Ref. [75] using the PARPACK library [98], from which we computed the corresponding unfolded eigenvalues. We complemented these data by performing a new calculation to obtain the IPRs of the Dirac eigenvectors for the gauge ensembles with $N_{t}=14$. We calculated error bars from a standard binned jack-knife performed over the whole analysis, including the unfolding procedure (when studying spectral statistics), to take into account possible correlations introduced by the latter. An example of a staggered Dirac spectrum on a single, typical configuration is shown in Fig. 1. We performed unfolding by ordering all the available eigenvalues $\lambda_{n}$ of all the configurations of a given lattice ensemble, and replacing them with their rank divided by the number of configurations [17]. This automatically yields unit spectral density, while the deviation of $\langle s\rangle$ from 1 can be used to test the accuracy of the procedure.
@@
Figure 2: Top panel: comparison of $\langle s\rangle$ as a function of $\lambda/m_{s}$ for all the explored lattice spacings at fixed value of the aspect ratio $LT=4$. Bottom panel: comparison of $\langle s\rangle$ as a function of $\lambda/m_{s}$ for all the explored aspect ratios, corresponding to different values of $N_{s}$, at fixed value of the lattice spacing, corresponding to a temporal extent $N_{t}=14$. Our continuum estimate for the mobility edge [see Eq. (44) in Sec. 3.6] is marked in both panels by a red circle on the horizontal axis.
@@ 3.3 Determination of the mobility edge from spectral statistics
Figure 4: Comparison of $I_{s_{0}}$ (top panel) and $\mathrm{Var}(s)$ (bottom panel), computed locally in the spectrum on lattices at fixed temperature $T=230\,\mathrm{MeV}$, for different values of the lattice spacing at fixed spatial lattice size $L=3.4\,\mathrm{fm}$ in physical units (i.e., fixed $LT=4$). Our continuum estimate for the mobility edge [see Eq. (44) in Sec. 3.6] is marked by a red circle on the horizontal axis.
Figure 5: Comparison of $I_{s_{0}}$ (top panel) and $\mathrm{Var}(s)$ (bottom panel), computed locally in the spectrum on lattices at fixed temperature $T=230\,\mathrm{MeV}$, for different spatial volumes at fixed lattice spacing, $a=0.0615\,\mathrm{fm}$ (i.e., $N_{t}=14$). Our continuum estimate for the mobility edge [see Eq. (44) in Sec. 3.6] is marked by a red circle on the horizontal axis.
@@
The data show that the two coarsest lattices ($N_{t}=8,10$) give compatible results for the spectral statistics in all spectral regions in the bulk of the spectrum, and down to the first bin below the mobility edge. Barring a conspiracy, this indicates two things: that the formation of taste multiplets does not have significantly large effects in that spectral region yet; and that, when this is the case, further lattice artifacts, on top of those introduced by taste-degenerate multiplets, are small when considering localization properties at fixed physical volume. Clearly, there could still be finite-volume systematics, but these would involve only effects unrelated to taste multiplets (such as the localization length in the spectral region near the mobility edge being too large compared with the spatial size of the system), which are expected not to affect much the determination of the mobility edge via the matching with the critical statistics (see above in Sec. 2.3). Indeed, previous numerical results show that for lattices of similar size and spacing one is already close to the one-parameter scaling regime near the mobility edge [5]. This shows that the determination of the mobility edge from the critical spectral statistics is reliable on these lattices.
The dependence of $\lambda_{c}^{\mathrm{stat}}/m_{s}$ on the aspect ratio at fixed lattice spacing is visible instead in Fig. 5, where $N_{t}=14$ and the spatial size is varied. Here a larger aspect ratio drives $I_{s_{0}}$ and $\mathrm{Var}(s)$ down, as expected, so leading to an estimate for the mobility edge that decreases with the spatial volume. Notice that the scale-invariant nature of the mobility edge is still masked here by the distortions of the unfolded level spacing distribution, and no volume-independent crossing point of the various curves is present. This shows that the effects of taste degeneracy are still strong even on our largest lattice at $N_{t}=14$. This is in agreement with the argument discussed in Sec. 2.4, as, according to Eq. (15), we would need an aspect ratio of $LT\sim 9$ at this lattice spacing in order to have the same $\delta/\Delta$ that we have for $N_{t}=8$ with $LT=4$. We then expect that the plateau in $\lambda_{c}^{\mathrm{stat}}/m_{s}$ against $LT$ is not visible yet, and would appear for an aspect ratio roughly twice as big as our largest one.
@@ 3.4 Taste-degeneracy effects on the mobility edge in the thermodynamic limit
In order to show that the effects of taste degeneracy become irrelevant for the determination of the mobility edge with our method when taking the thermodynamic limit at fixed lattice spacing, we have compared the $N_{s}\to\infty$ extrapolation of our $N_{t}=14$ data with an independent determination of $\lambda_{c}$ based directly on the properties of the eigenvectors. Distortive effects on the statistics are due to approximate taste degeneracy, which is visible when $(a\Lambda)^{2}<1/(V\Lambda^{3})$. We then expect finite-size corrections to $\lambda_{c}^{\mathrm{stat}}/m_{s}$ of order $O(1/N_{s}^{3})$, up to lattice sizes where the effects of taste degeneracy disappear. As these are expected to be about an order of magnitude larger than our largest lattices, we can ignore the onset of the plateau in our extrapolation in $(N_{t}/N_{s})^{3}$. The infinite-volume extrapolation is shown in Fig. 7; the corresponding results for the mobility edge are:
@@
This is in good agreement with the thermodynamic extrapolation of our estimates from spectral statistics, Eq. (37), see Fig. 7. This shows that these are indeed converging to the actual position of the mobility edge, as we argued in Sec. 2.4. Moreover, this further confirms that finite-size effects on the estimate based on the critical fractal dimension are reasonably small.
@@
3.5 Dependence of the mobility edge on the lattice spacing and exchange of continuum and thermodynamic limits
@@
We finally summarize the results of the previous two subsections in Fig. 10, showing together all our determinations of $\lambda_{c}/m_{s}$. It is clear that our two coarsest lattices give compatible estimates via $\lambda_{c}^{\mathrm{stat}}/m_{s}$, while on finer lattices with the same aspect ratio these estimates rapidly deviate. On the other hand, the infinite-volume extrapolation at $N_{t}=14$ gives again results compatible with our coarsest lattices, as well as with the determination from the local fractal dimension.
@@
Before attempting an extrapolation of $\lambda_{c}/m_{s}$ to the continuum, it is important to check that our numerical results are in agreement with our theoretical argument developed in Sec. 2.5 that the thermodynamic and continuum limits can be exchanged. While a direct check is not possible with the available data, indirect support is provided by our results for the estimate based on $\alpha(\lambda)$. Being unaffected by taste degeneracy, this gives us a valid estimate of the position of the mobility edge at both finite aspect ratio and finite spacing. We have already pointed out above that this direct estimate of the mobility edge does not depend strongly on the aspect ratios employed, and agrees with the indirect estimate from spectral statistics extrapolated to infinite volume. Moreover, our results obtained at $N_{t}=14$ are not very different from the estimates obtained using spectral statistics on our coarsest lattices $N_{t}=8,10$, where taste-degeneracy effects on spectral statistics are under control and estimates based on critical statistics should accurately reflect the position of the mobility edge (within our numerical accuracy). These observations suggest that the true position of the mobility edge, as intrinsically defined by the properties of the eigenvectors, does not depend strongly neither on the aspect ratios nor on the lattice spacing. This is in agreement with our argument that the continuum and thermodynamic limit commute for what concerns the mobility edge. We can then use indirect estimates of the mobility edge from spectral statistics extrapolated to the thermodynamic limit, followed by the continuum limit, to obtain the correct value of the mobility edge in the continuum theory.
Further numerical support to the correctness of our theoretical argument can be obtained if we manage to remove the effects of taste degeneracy from the spectrum. For a suitably “taste-symmetrized” spectrum, one expects a more accurate correspondence between localization properties of the eigenvectors and spectral statistics already at finite volume. Agreement of the resulting estimates for the mobility edge with the ones obtained by using the spectrum without further processing and taking the thermodynamic limit, would provide further support to the correctness of our procedure. Clearly, the taste-symmetrization of the spectrum is an ambiguous procedure, even if one could correctly identify the eigenvalue multiplets, which is not easy when the spacing is too coarse; even on our finest lattices, only doublets are visible. This ambiguity leads to hardly quantifiable systematic effects, that make this approach less reliable for a controlled continuum extrapolation.
To obtain a heuristic estimate we have then opted for the simplest approach, namely replacing pairs of neighboring eigenvalues with their arithmetic average. This has a number of obvious theoretical issues, first of all the likely misidentification of at least part of the doublets. This problem is particularly severe in the bulk of the spectrum, where doublet misidentification is almost certain, leading to interfering with the relevant spectral correlations and possibly to destroying them completely. Nonetheless, one expects that for finer and finer lattices at fixed aspect ratio this procedure should work better and better (and worse and worse on bigger and bigger lattices at fixed lattice spacing), and eventually provide the correct result in the continuum limit (if one extends it to quartets when these show up); and one can hope that for sufficiently fine but still manageable lattices it works reasonably well up to and slightly above the mobility edge. We have then averaged the spectrum as discussed above, and performed on it the same analysis as with the untreated spectrum, looking for the point where spectral statistics match their critical value. We restricted the spectrum to regions where the unfolded symmetrized level spacings satisfy $\langle s\rangle\simeq 1$, as a minimum requirement for the reliability of the procedure. This is illustrated in Fig. 11.
With this long list of caveats in mind, we show our results in Figs. 11 and 12. As expected, RMT correlations in the bulk of the spectrum are lost. However, in spite of our rather crude approach, the agreement between estimates based on the taste-averaged and untreated spectra is surprisingly good already for our $N_{t}=14$ lattices. While uncertainties are hard to quantify, we believe that these results provide strong numerical support to our theoretical argument for the exchangeability of continuum and thermodynamic limits.
For future utility, we report also the results obtained with the taste-averaging procedure on our finest, $N_{t}=16$ lattice, i.e., $\lambda_{c}/m_{s}=1.754(33)$ from $I_{s_{0}}$ and $\lambda_{c}/m_{s}=1.766(54)$ from $\mathrm{Var}(s)$.
Figure 11: Determination of the mobility edge from the spectral statistics of the spectrum after taste averaging (see text for details), estimated using results from $48^{3}\times 14$ lattices.
Figure 12: Estimates for the mobility edge from the spectral statistics of the spectrum after taste averaging (see text for details), obtained using results from our $N_{t}=14,16$ lattices. For comparison, we show also our results from the fractal dimension and from standard spectral statistics extrapolated to the thermodynamic limit on $N_{t}=14$ lattices, and our continuum extrapolation based on standard spectral statistics (see Sec. 3.6).
3.6 Continuum extrapolation
@@
In the light of this discussion, we have performed an extrapolation to the continuum limit using our estimates of $\lambda_{c}/m_{s}$ from spectral statistics for $N_{t}=8,10$ and $N_{t}=14$, where for $N_{t}=14$ we have used the results extrapolated first toward the thermodynamic limit. For the continuum extrapolation we have assumed the behavior $\frac{\lambda_{c}}{m_{s}}=\frac{\lambda_{c}}{m_{s}}\big{|}_{\mathrm{cont}}+C/N_{t}^{2}$, based on our theoretical expectations discussed in Sec. 2.5. The results obtained from the two observables $I_{s_{0}}$ and $\mathrm{Var}(s)$ are:
@@
$\frac{\lambda_{c}}{m_{s}}\bigg{|}_{\mathrm{cont}}$ $=1.748(75),\qquad$ (from $(\text{from }I_{s_{0}}),$ (41)
$\frac{\lambda_{c}}{m_{s}}\bigg{|}_{\mathrm{cont}}$ $=1.816(43),\qquad$ (from $(\text{from }\mathrm{Var}(s)),$ (42)
perfectly compatible within errors. We note that, while affected by uncontrolled theoretical uncertainties, the results obtained with our taste-averaging procedure on our finest, $N_{t}=16$ lattice, reported above at the end of Sec. 3.5, are in excellent agreement with these estimates (see Fig. 12).
To quote a final number, we did a weighted average of our two estimates and their statistical errors, and used their difference as an estimate of systematic finite-size effects. We find:
$\frac{\lambda_{c}}{m_{s}}\bigg{|}_{\mathrm{cont}}$ $=1.799(51)_{\mathrm{stat}}(68)_{\mathrm{syst}},$ (43)
$=1.799(85)_{\mathrm{combined}}.$ (44)
@@
To our knowledge, this result is the first fully controlled extrapolation of $\lambda_{c}/m_{s}$ to the continuum.